# Conan: A Chunkwise Online Network for Zero-Shot Adaptive Voice Conversion

### v1

@@ I Introduction
The field of VC has seen substantial advances [4]. In particular, zero‐shot VC, which aims to transfer out-of-domain speaker identity from reference speech to generated speech, has been explored in recent years [5]. Moreover, real‐time VC systems capable of operating online have been successfully deployed in low‐latency scenarios [6]. However, with the recent growing demand for customizable, high‐quality audio experiences, there is an emerging trend toward zero‐shot online VC that delivers high semantic fidelity, adaptive style transfer, and high-level naturalness [1]. For chunkwise online source speech input, it is essential not only to maintain both content extraction and speech synthesis quality under strict online latency constraints but also to transfer the timbre and styles of unseen reference speakers. Such speech styles encompass emotion, prosody, and other expressive attributes [7], adapting these out‐of‐domain timbre and styles to arbitrary source content in VC remains a challenging open problem.
@@
  • Achieving adaptive style transfer in zero-shot scenarios is still challenging. Current models often rely on the strong assumption that the reference voice is accessible [11] for model adaptation, which does not always hold in practice. Moreover, many VC approaches focus solely on timbre transfer by directly applying the source speech’s F0 and energy, which contradicts the goal of capturing the reference speaker’s styles [3]. The handful of zero-shot VC methods that do attempt to generalize to unseen speakers still struggle to balance high speaker similarity with naturalness. This shortfall arises from insufficient modeling of speaker styles, preventing effective alignment of speaker styles with source content [6].
@@ III-B Stream Content Extractor
In this attention mechanism, keys, values, and memory‐bank vectors are employed to capture contextual dependencies between data chunks. By flexibly adjusting the size of the right context, one can trade off performance against latency, setting it to zero enforces a strictly causal behavior. The sizes of the left context and memory buffer determine the overall receptive field. Leveraging our Stream Content Extractor, we achieve efficient, high‐quality streaming content extraction.
@@ IV-A Experimental Setup
Evaluation Details. We objectively and subjectively evaluate our model based on three criteria: real‐time content accuracy, conversion quality, and speaker similarity. Content accuracy is measured by Word Error Rate (WER) and Character Error Rate (CER) using a HuBERT‐Large ASR model  [^1] https://github.com/resemble-ai/Resemblyzer. Speaker similarity (SIM) is assessed with Resemblyzer  [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. For subjective metrics, we conduct mean opinion score (MOS) and comparative mean opinion score (CMOS) evaluations. Both metrics are rated on a 1–5 scale and reported with 95% confidence intervals. Specifically, we use MOS‐Q to judge conversion quality (for clarity and naturalness) and MOS‐S to assess speaker similarity (for timbre and styles). For CMOS-Q and CMOS-S, listeners are tasked with comparing pairs of speech produced by different systems and expressing their preferences. All participants are fairly compensated and agree that the results will be used for scientific research.

### v2

@@ I Introduction
The field of VC has seen substantial advances [4]. In particular, zero‐shot VC, which aims to transfer out-of-domain speaker identity from reference speech to generated speech, has been explored in recent years [5]. Moreover, real‐time VC systems capable of operating online have been successfully deployed in low‐latency scenarios [6]. However, with the recent growing demand for customizable, high‐quality audio experiences, there is an emerging trend toward zero‐shot online VC that delivers high semantic fidelity, adaptive style transfer, and high-level naturalness [1]. For chunkwise online source speech input, it is essential not only to maintain both content extraction and speech synthesis quality under strict online latency constraints but also to transfer the timbre and styles of unseen reference speakers. Such speech styles encompass emotion, prosody, and other expressive attributes [7]. Adapting these out‐of‐domain timbres and styles to arbitrary source content in VC remains a challenging open problem.
@@
  • Achieving adaptive style transfer in zero-shot scenarios is still challenging. Current models often rely on the strong assumption that the reference voice is accessible [11] for model adaptation, which does not always hold in practice. Moreover, many VC approaches focus solely on timbre transfer by directly applying the source speech’s F0 and energy, which contradicts the goal of capturing the reference speaker’s styles [3]. The handful of zero-shot methods that do attempt to generalize to unseen speakers still struggle to balance high speaker similarity with naturalness [12]. This shortfall arises from insufficient modeling of speaker styles, preventing effective alignment of speaker styles with source content [6].
@@ III-B Stream Content Extractor
In this attention mechanism, keys, values, and memory‐bank vectors are employed to capture contextual dependencies between data chunks. By flexibly adjusting the size of the right context, one can trade off performance against latency. Setting it to zero enforces a strictly causal behavior. The sizes of the left context and memory buffer determine the overall receptive field. Leveraging our Stream Content Extractor, we achieve efficient, high‐quality streaming content extraction.
@@ IV-A Experimental Setup
Evaluation Details. We objectively and subjectively evaluate our model based on three criteria: real‐time content accuracy, conversion quality, and speaker similarity. Content accuracy is measured by Word Error Rate (WER) and Character Error Rate (CER) using a HuBERT‐Large ASR model  [^1] https://github.com/resemble-ai/Resemblyzer. Speaker similarity (SIM) is assessed with Resemblyzer  [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. For subjective metrics, we conduct mean opinion score (MOS) and comparative mean opinion score (CMOS) evaluations. Both metrics are rated on a 1–5 scale and reported with 95% confidence intervals. Specifically, we use MOS‐Q to judge conversion quality (for clarity and naturalness) and MOS‐S to assess speaker similarity (for timbre and styles) [47]. For CMOS-Q and CMOS-S, listeners are tasked with comparing pairs of speech produced by different systems and expressing their preferences. All participants are fairly compensated and agree that the results will be used for scientific research.